OpenStudy (anonymous):

2. What is the distance between P(–4, 3) and Q(6, 1)? Round to the nearest tenth. (2 points) PLEASE EXPLAIN

OpenStudy (austinl):

\(d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\)

OpenStudy (anonymous):

So D+ (6 -(-4) )^2 + (1-3)^2 ?

OpenStudy (anonymous):

*D=

OpenStudy (austinl):

Yep, and square root of it too. \(d=\sqrt{(6-(-4))^2+(1-3)^2}=~\sqrt{10^2+(-2)^2}=~?\)

OpenStudy (austinl):

@wealdooo That is the slope formula....... and they are subscripts not exponents in that particular formula.

OpenStudy (anonymous):

Soooo i got 10.19803

OpenStudy (austinl):

Correct, but incorrect rounding. What is it to the nearest tenth?

OpenStudy (anonymous):

10.2??

OpenStudy (austinl):

Correctomundo!

OpenStudy (anonymous):

yay thank you so much!

OpenStudy (austinl):

No problem!